What is the molar concentration of a solution that has 5.00g of ammonium chloride dissolved in 250.0mL of solution 


1
Expert's answer
2022-01-27T00:04:01-0500

Answer - Given, mass of ammonium chloride = 5.00 g

Volume = 250.0 mL

Step 1) Moles of ammonium chloride-

Mole = mass ÷ molar mass.

Moles of NH4Cl ="\frac{5.00 g}{53.49} g/mol"  = 0.0935 moles.

Step 2) Molar Concentration -

We know formula,

Molarity ="\frac{moles}{volume (L)}"

Volume = 250 mL = 0.250 L

Molarity of NH4Cl ="\frac{0.0935 moles}{ 0.250 L}" = 0.374 M